# Safe Reinforcement Learning In Black-Box Environments Via Adaptive Shielding

### v2

@@ Appendix D Parameter Analysis
Figure 12: The rate of safety violations, and rate of goal reaches when ADVICE is activated at different intervals $E$.
ADVICE has a cold-start, meaning it requires some period of time before activation to collect data in order to work efficiently. We acknowledge that this can affect the performance of ADVICE  evidently we show the results of various activation points in Figure 12. To allow for a fair comparison as possible, we show the rate of safety violations for the subsequent $1000$ episodes after activation. Again, we visualise a trade-off. Delaying ADVICE’s activation for longer results in fewer safety violations and increased goal reaches. However, the RL agent observes more cumulative safety violations up to the point of activation. Starting ADVICE earlier decreases the number of safety violations up to activation but gives the autoencoder fewer data points to train off. Evidently, safety violations are not reduced to the same magnitude and goal reaches also decrease. This is to be expected with any neural network-based approach.
Figure 13: Average episodic reward, cumulative safety violations, and cumulative goal reaches of various methods on the randomised goal environment where the maximum episodic steps are doubled.
Based on results in Figure 2(b), we hypothesise that the reduced reward and cumulative goal reaches is a result of ADVICE not having enough time to complete the task. As seen in Figure 3, ADVICE learns to take a longer route to the objective, so by doubling the maximum step count allowed per episode, we expect to see an increase in cumulative goal reaches, average reward, and no increase to safety violations. Results for this experiment are shown in Figure 13. As expected, given more time to complete the task, ADVICE now reaches the goal more than when the maximum step counter is the default value. As a result, we observe an increase in average episodic return much closer to baseline methods.
@@ Appendix E Adaptive ADVICE
Consecutive Episodes ($K=4$) $22.60\pm 6.71$ $22.10\pm 7.26$
Consecutive Episodes ($K=5$) $00.00\pm 0$ 0.00 $181.35\pm 4.92$
@@ Appendix F Hyperparameter Analysis and Computational Overheads
This section lists the hyperparameters used by all models and ADVICE. Table 2 summarises all hyperparameters used in Section 6. We will refer the reader to our source code repository 1 for the remaining details.
Using all model configurations in Table 2, a single ADVICE run (one random seed) takes 12, 24, 24, 12, and 12 hours of training, respectively, in the semi-random goal, random goal, random button, random circle, and constrained random goal environments. For all other methods, a single run takes 3, 4, 5, 3, and 5 hours in the same environments, respectively. All experiments were run on a large computing cluster utilising two Nvidia H100 GPUs, 16 CPUs, and up to 500GB memory.
@@
Batch size 64 No. Neighbours ($K^{max}$) 5
@@
Hyperparameters for the DDPG algorithm started with author recommendations [Lillicrap et al., 2015]. They were manually tuned afterwards to achieve a high performance on individual environments before tests were carried out, meaning the RL algorithm for all approaches was of high performance and fair comparison. Hyperparameters for ADVICE were manually tuned for performance in CL loss and MSE loss. Some hyperparameter analysis was conducted in Section D to justify certain choices. Parameters for the DDPG-Lag method started with recommendations [Stooke et al., 2020] and were tuned for performance in our experiments.

### v3

@@ Appendix D Parameter Analysis
Figure 14: The rate of safety violations, and the rate of goal reaches when ADVICE is activated at different intervals $E$.
ADVICE has a cold-start, meaning it requires some period of time before activation to collect data in order to work efficiently. We acknowledge that this can affect the performance of ADVICE  evidently, we show the results of various activation points in Figure 14. To allow for a fair comparison as possible, we show the rate of safety violations for the subsequent $1000$ episodes after activation. Again, we visualise a trade-off. Delaying ADVICE’s activation for longer results in fewer safety violations and increased goal reaches. However, the RL agent observes more cumulative safety violations up to the point of activation. Starting ADVICE earlier decreases the number of safety violations up to activation, but gives the autoencoder fewer data points to train on. Evidently, safety violations are not reduced to the same magnitude, and goal reaches also decrease. This is to be expected with any neural network-based approach.
Figure 15: Average episodic reward, cumulative safety violations, and cumulative goal reaches of various approaches on the randomised goal environment where the maximum episodic steps are doubled.
Based on results in Figure 2b, we hypothesise that the reduced reward and cumulative goal reaches is a result of ADVICE not having enough time to complete the task. As seen in Figure 3, ADVICE learns to take a longer route to the objective, so by doubling the maximum step count allowed per episode, we expect to see an increase in cumulative goal reaches, average reward, and no increase in safety violations. Results for this experiment are shown in Figure 4. As expected, given more time to complete the task, ADVICE now reaches the goal more than when the maximum step counter is the default value. As a result, we observe an increase in average episodic return much closer to baseline approaches.
@@ Appendix E Adaptive ADVICE
Consecutive Episodes ($K=4$) $2.60\pm 6.71 22$$2.10\pm 7.26$
Consecutive Episodes ($K=5$) $0.00\pm 0.00$ $181.35\pm 4.92$
@@ Appendix F Hyperparameter Analysis and Computational Overheads
This section lists the hyperparameters used by all models and ADVICE. Table 2 summarises all hyperparameters used in Section 6. We will refer the reader to our source code repository for the remaining details.
Using all model configurations in Table 2, a single ADVICE run (one random seed) takes 12, 24, 24, 12, and 12 hours of training, respectively, in the semi-random goal, random goal, random button, random circle, and constrained random goal environments. For all other approaches, a single run takes 3, 4, 5, 3, and 5 hours in the same environments, respectively. All experiments were run on a large computing cluster utilising two Nvidia H100 GPUs, 16 CPUs, and up to 500GB memory.
@@
Batch size 64 No. Neighbours ($K_{max}$) 5
@@
Hyperparameters for the DDPG algorithm started with author recommendations [29]. They were manually tuned afterwards to achieve a high performance on individual environments before tests were carried out, meaning the RL algorithm for all approaches was of high performance and fair comparison. Hyperparameters for ADVICE were manually tuned for performance in CL loss and MSE loss. Some hyperparameter analysis was conducted in Section D to justify certain choices. Parameters for the DDPG-Lag approach started with recommendations [36] and were tuned for performance in our experiments.